OpenStudy (charlest):

Simplify. 16 – 7 – (–4) + 12 A. –7 B. 17 C. 25 D. 39

OpenStudy (janu16):

C 25

OpenStudy (charlest):

When I do the equation though I get 17.

OpenStudy (janu16):

simplify brackets 16−7+4+12 16−7+4+12 2 Simplify 25

OpenStudy (janu16):

16 -7 is 9 than 9 - -4 is 13 than 13 +12 is 25

OpenStudy (janu16):

you did something wrong

OpenStudy (charlest):

16-7=9 9-4=5 5+12=17 I keep getting 17

OpenStudy (janu16):

its 9- negative 4

OpenStudy (janu16):

do you see it?

OpenStudy (janu16):

you did 9 - 4 but its supposed to be 9 - negative 4

OpenStudy (charlest):

I'm subtracting the absolute values because they're not the same sign

OpenStudy (janu16):

if you pick 17, its gonna be wrong

OpenStudy (janu16):

ok look let me explain

OpenStudy (janu16):

what is 16 -7?

OpenStudy (charlest):

9

OpenStudy (charlest):

9-(-4)= 5

OpenStudy (janu16):

no its 13

OpenStudy (charlest):

Im supposed to subtract the absolute values if the signs are different

OpenStudy (janu16):

can you put that in calcultaor

OpenStudy (janu16):

they are not different

OpenStudy (janu16):

look if you put 9-(-4) in calc you will get 13 than +12 is 25

OpenStudy (charlest):

no 9 is positive and 4 is negative that's what im not getting

OpenStudy (janu16):

you are doing 9 MINUS(negative 4)

OpenStudy (janu16):

even though 9 is positvie, its telling you to subtract

OpenStudy (charlest):

wait, why are you adding?

OpenStudy (janu16):

im not check thsi link plz

OpenStudy (janu16):

check this and tell me if you understand

OpenStudy (charlest):

oh. It's turning the subtraction sign into a addition sign. ok i get it now

OpenStudy (charlest):

so it's C?

OpenStudy (janu16):

no i see what you did i think

OpenStudy (janu16):

nvm

OpenStudy (janu16):

you understand right?

OpenStudy (charlest):

In my lesson, My teacher told me to subtract the absolute values if the signs are different

OpenStudy (janu16):

ok i dont know what that means cause i didnt learn tht but do you understand what im showing?

OpenStudy (charlest):

im subtracting the absolute values instead of adding them.

OpenStudy (janu16):

ohok

OpenStudy (janu16):

so its C 25

OpenStudy (charlest):

yeah

OpenStudy (charlest):

ok I'll see what i get

OpenStudy (janu16):

yup

OpenStudy (janu16):

let me know cause im 100 percent sure its correct and you even saw the link

OpenStudy (charlest):

40% (TT_TT)

OpenStudy (charlest):

I think i was right

OpenStudy (janu16):

ok what did you pcik

OpenStudy (jjcoolja12):

16−7−(−4)+12 =25

OpenStudy (charlest):

no you were right